# Relativistic fluid dynamics in a ‘hydro’ frame

### v1

@@ Abstract
In relativistic hydrodynamics, defining fluid velocity and temperature beyond the perfect fluid case introduces ambiguity. In thermodynamic equilibrium, these variables are naturally defined by uniform thermodynamic potentials. Outside equilibrium, they can be redefined with arbitrary corrections that become nonzero with spacetime variations. These ambiguities are typically resolved by imposing constraints on the conserved currents, akin to gauge fixing in gauge theory, with common choices being the ‘Landau frame’ or ‘Eckart frame’ [7, 8, 9]. However, one can always opt for constraints beyond these two options or even proceed without addressing these field ambiguities, as demonstrated in BDNK (Bemfica-Disconzi-Noronha-Kovtun) formalism [10, 11]. It appears that the spectrum of linearized perturbations is significantly altered when the fluid fields are redefined. The BDNK formalism, in particular, leaves the fluid frame unfixed to establish the frame in which the first-order fluid dynamics preserves causality and stability. This method diverges from the established difficulties related to causality and stability in the traditional Landau or Eckart frame [12, 13, 14, 15, 16, 17]. While rotational and translational symmetries are easily addressed in the analysis, gauge freedoms in defining fluid variables add further complexities. Addressing these complexities is crucial for ensuring that the physical constraints derived from linearized perturbation analysis appropriately represent the underlying physics without any artefacts from arbitrary definitions [18, 19, 20, 21, 22, 23]. It is well understood that the linearized spectrum changes with field redefinition, as the perturbations of the redefined fields correspond to different physical quantities than the original fields [24, 25, 26]. However, the causal nature of a specific fluid model should not be contingent upon the choice of variables used for its description.
@@
We, in this work, provided a novel technique to find out which fluid model is physical under arbitrary field redefinition. The method is simple: we, through an all-order field redefinition, eliminate all non-hydrodynamic modes from the fluid model and analyze the structure of the resulting modified energy-momentum tensor. This technique works provided the original stress-energy tensor (the one with non-hydrodynamic modes, before the field redefinition) has a finite number of terms in derivative expansion. We observe that if the new modified stress-energy tensor, after eliminating the non-hydrodynamic modes, incorporates higher derivative terms of the fluid variables that include up to infinite orders, then the fluid theory we started with is physical, whereas, if the modified stress-energy tensor turns out to have a finite number of terms, then the non-hydrodynamic modes of the original theory must emerge from field redefinition artefacts solely and should not be considered physical. Another criterion we find is whether the dispersion polynomial neatly factorizes into hydrodynamic and non-hydrodynamic modes. If the non-hydrodynamic modes do not appear as distinct finite factors, then they are not frame artefacts [28].
In this and a companion study [28], we aim to investigate this impact of field redefinition on the spectrum of linearized perturbations within relativistic fluid dynamics. We have found that the spectrum of hydrodynamic modes (those with frequencies that vanish as spatial momentum approaches zero) remains unchanged by local field redefinition. In contrast, the spectrum of non-hydrodynamic modes (those with frequencies that approach a finite nonzero value as momenta diminish) is influenced by field redefinition. We also show that non-hydrodynamic modes can be removed from the spectrum of linearized perturbations through a suitable, generally all-order field redefinition. Once this is applied, the spectrum will be simplified to include only the hydrodynamic modes of the original fluid model, and the stress-energy tensor becomes ready for analyzing the physical validity of the original non-hydrodynamic modes. We will refer to this new fluid frame as the “hydro frame”. Even if we start with a stress tensor that contains a finite number of terms, transforming to the “hydro frame” typically yields a stress tensor with an infinite number of terms, incorporating arbitrarily higher-order derivatives of the fluid variables. In Ref. [28], we have shown and proven how the non-hydrodynamic modes of the original theory could control the hydrodynamic expansion’s validity of the new stress-energy tensor.
@@
where $M_{ij}$ is the linearization matrix, and the spectrum is obtained from the zeros of the determinant of $M_{ij}$. If the equation ${\cal E}$ contains a finite number of derivatives, then $\text{Det}[M]$ will be a finite polynomial in $\omega$ and $k$, resulting in a finite number of zeros. Now consider a field redefinition $\Phi_{i}\rightarrow\Psi_{i}=\Phi_{i}+\Delta\Phi_{i}$, where the shift $\Delta\Phi_{i}$ is a nonlinear function of $\Phi_{i}$ and its derivatives. In the context of fluid dynamics, we assume that at equilibrium, $\Phi_{i}$ and $\Psi_{i}$ coincide, meaning $\Delta\Phi_{i}$ vanishes when evaluated at $\Phi_{i}=\bar{\Phi}_{i}$. Under this field redefinition, the equations of motion transform as ${\cal E}(\Phi)\rightarrow\tilde{\cal E}(\Psi)$. While the set of PDEs $\tilde{\cal E}$ can be fully determined from ${\cal E}$ and the transformation $\Delta\Phi$, it will have a distinctly different structure. We can linearize $\tilde{\cal E}$ to obtain the spectrum, similarly to how we did for ${\cal E}$ in (1), resulting in a different linearization matrix $\tilde{M}_{ij}$. Our objective is to establish the relationship between the matrices $M_{ij}$ and $\tilde{M}_{ij}$
@@
As $\Delta\Phi$ is vanishing when $\Phi=\bar{\Phi}$, it must be of order $\mathcal{O}(\epsilon)$ or higher for the field configuration in (1). Thus, $\Delta\Phi$ can generally be represented by the matrix equation up to ${\cal O}(\epsilon^{2})$
@@
$\Rightarrow Det[M]=Det[\tilde{M}]Det[{\bf 1}+S].$
The zeros of ${Det}[{\bf 1}+S]$ represent new modes in the $\Phi$ frame that are absent in the $\Psi$ frame. These new zero modes arise from perturbations in the $\Phi$ frame, where the additional terms coming from the linearized field redefinitions are exactly canceled by the linearized fluctuations. In the $\Psi$ frame, these new modes indicate no fluctuations, meaning that they are artefacts of the frame transformation with no physical significance. Importantly, depending on the choice of field redefinitions, these artificial modes may appear unstable or acausal, even if the theory in the $\Psi$ frame is entirely valid. However, in the $\Phi$ frame, there is no straightforward way to distinguish these artefacts from genuine physical modes.
@@
$\begin{split}F_{u(T)}&\equiv\sum_{m,n}f_{m,n}^{u(T)}\left[\hat{u}\cdot\partial\right]^{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{n}\,,\\ G_{u(T)}&\equiv\sum_{m,n}g^{u(T)}_{m,n}\left[\hat{u}\cdot\partial\right]^{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{n}\,,\end{split}$ (9)
and $R_{u(T)}=R_{u(T)}[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}]$ do not depend on the operator $(\hat{u}\cdot\partial)$. They are included to account for field redefinitions that remain non-zero as $\omega\rightarrow 0$. These can also be expanded in the small $k$ limit as $R_{u(T)}=\sum_{m}r^{u(T)}_{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{m}$. Here, $f_{m,n}^{u(T)}$, $g^{u(T)}_{m,n}$, and $r^{u(T)}_{m}$ are functions of temperature without its derivatives, and $\hat{P}^{\mu\nu}$ is the projector orthogonal to $\hat{u}^{\mu}$, $\hat{P}^{\mu\nu}=\eta^{\mu\nu}+\hat{u}^{\mu}\hat{u}^{\nu}$. In the context of linearized analysis, the derivatives $(\hat{u}^{\alpha}\partial_{\alpha})$ and $(\hat{P}^{\alpha\beta}\partial_{\alpha})$ commute, so their ordering in the functions don’t matter. For a rotational and translationally invariant equilibrium case, we have $\bar{u}^{\mu}=\{1,0,0,0\},\quad\bar{T}={\rm constant}$, where the perturbed solution is $u_{\mu}^{(s)}=\bar{u}_{\mu}+\epsilon~{}\delta u_{\mu}\,e^{-i\omega t+i\vec{k}\cdot\vec{x}},\quad T^{(s)}=\bar{T}+\epsilon\,\delta T\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$.
@@
where $\Delta u_{k}$ and $\Delta u_{\perp}$ are defined the same way as $\beta_{k}$ and $\beta_{\perp}$. Now we compute the determinant of $\left[{\mathbb{1}}+S\right]$ mentioned in Eq. (6) as
${\cal F}(-i\omega,k^{2})\equiv Det[{\mathbb{1}}+S]=(1-i\omega F_{u}-k^{2}R_{u})\times$
$\left[(1-i\omega G_{T}-k^{2}R_{T})(1-i\omega F_{u}-k^{2}R_{u})+k^{2}F_{T}~{}G_{u}\right].\quad$ (12)
We have suppressed the arguments of $F_{u(T)}$, $G_{u(T)}$, and $R_{u(T)}$ for simplicity. From Eq. (6), we note that a general field redefinition of the fluid variables introduces an additional factor, $\mathcal{F}$, to the dispersion polynomial which is a polynomial in $\omega$ and $k^{2}$. This generates new modes in the system.
Some remarks on the properties of $F_{u(T)}$ and $G_{u(T)}$: If $F_{u}$ and $G_{T}$ are vanishing, and $F_{T}$ and $G_{u}$ do not depend on the operator $(u\cdot\partial)$ or $(-i\omega)$ in Fourier space, then $\mathcal{F}$ will have no zeros for any $\omega$, resulting in no new modes. However, depending on $F_{T}$ and $G_{u}$, $\mathcal{F}$ might have zeros for some real values of $k$. If every term in $F_{u}$ and $G_{T}$ includes at least one factor of $(P^{\mu\nu}\partial_{\mu}\partial_{\nu})$ or $k^{2}$ in Fourier space, the frequency of these new modes will diverge as $k\rightarrow 0$. If either $F_{u}$ or $G_{T}$ has at least one term without $(P^{\mu\nu}\partial_{\mu}\partial_{\nu})$, at least one new mode will appear with a finite, non-zero frequency as $k\rightarrow 0$, resembling a genuine “non-hydrodynamic” mode. None of the zeros of $\mathcal{F}$ will have the form $\lim_{k\rightarrow 0}\omega(k)=0$. Thus, frame redefinitions will not generate new hydrodynamic modes, and redefining fluid variables will not affect the spectrum of hydrodynamic modes in Fourier space.
@@
Now, let’s reverse this process—starting with the ‘Frame-2’ equations (using the hatted fluid variables $\hat{u}^{\mu}$ and $\hat{T}$) and implementing the inverse transformations from (7) and (8) to return to ‘Frame-1’ equation with fluid variables $u^{\mu}$ and $T$. This inversion would naturally eliminate the extra factor $\mathcal{F}$ from the ‘Frame-2’ dispersion polynomial. From the viewpoint of ‘Frame-2’, this effectively removes a non-hydrodynamic mode from the spectrum through field redefinition.
The possibility to fully absorb a non-hydrodynamic mode through a field redefinition implies that such a mode may not be physical. A truly physical non-hydrodynamic mode should never be completely erased by a frame transformation. In this section, we will investigate an “inverse transformation” to remove a mode from the spectrum and analyze if physical non-hydrodynamic modes persist as expected. The dispersion polynomials in ‘Frame-1’ and ‘Frame-2’, indicated as $P(\omega,k^{2})$ and $\hat{P}(\omega,k^{2})$, respectively, are related as
@@
$\text{where}\qquad{{\cal F}(\omega,k^{2})}\equiv(1-i\omega F_{u}-k^{2}R_{u})\times$
$\left[(1-i\omega G_{T}-k^{2}R_{T})(1-i\omega F_{u}-k^{2}R_{u})+k^{2}F_{T}~{}G_{u}\right]$
$\text{with}\quad F_{u(T)}$$\equiv$ $\sum_{m,n}f_{m,n}^{u(T)}\left[-i\omega\right]^{m}\left[-k^{2}\right]^{n},$
@@
Due to the isotropy of the background fluid profile and the constitutive relations, the linearized dynamics in the shear and sound sectors will get disentangled. Consequently, both $\hat{P}(\omega,k^{2})$ and $P(\omega,k^{2})$ must factor into contributions from the shear and sound sectors. The same holds for $\mathcal{F}(\omega,k^{2})$. To put it differently, we can separately recognize the factors as
@@
where $\hat{P}(\omega,k^{2})=\hat{P}_{\rm sh}(\omega,k^{2})\hat{P}_{{\rm snd}}(\omega,k^{2})$, and $P(\omega,k^{2})=P_{\rm sh}(\omega,k^{2})P_{{\rm snd}}(\omega,k^{2})$. Assume that $P_{\rm sh}(\omega,k^{2})$ has $N^{\rm sh}_{1}$ hydrodynamic modes with frequencies $\omega_{\rm sh}^{a}(k)$ and $N^{\rm sh}_{2}$ non-hydrodynamic modes with frequencies $\mathfrak{w}_{\rm sh}^{a}(k^{2})$ where, respectively, $a=\{1,2,\ldots,N^{\rm sh}_{1}\}$ and $a=\{1,2,\ldots,N^{\rm sh}_{2}\}$. And in the sound channel, it has $N^{\rm snd}_{1}$ hydrodynamic modes with frequencies $\omega_{\rm snd}^{a}(k)$ $\left(a=\{1,2,\ldots,N^{\rm snd}_{1}\}\right)$, and $N^{\rm snd}_{2}$ non-hydrodynamic modes with frequencies $\mathfrak{w}_{\rm snd}^{a}(k^{2})$ $\left(a=\{1,2,\ldots,N^{\rm snd}_{2}\}\right)$. Thus, $P^{\rm sh}(\omega,k^{2})$ and $P^{\rm snd}(\omega,k^{2})$ can be factored as $P^{\rm sh}(\omega,k^{2})=\left[\prod_{a=1}^{N^{\rm sh}_{1}}(\omega-\omega_{\rm sh}^{(a)})\right]\left[\prod_{a=1}^{N^{\rm sh}_{2}}(\omega-{\mathfrak{w}}_{\rm sh}^{(a)})\right]$, and $P^{{\rm snd}}(\omega,k^{2})=\left[\prod_{a=1}^{N^{{\rm snd}}_{1}}(\omega-\omega_{{\rm snd}}^{(a)})\right]\left[\prod_{a=1}^{N^{{\rm{\rm snd}}}_{2}}(\omega-{\mathfrak{w}}_{{\rm{\rm snd}}}^{(a)})\right]$. Both $\omega^{(a)}_{{\rm sh}/{\rm snd}}$ and $\mathfrak{w}^{(a)}_{{\rm sh}/{\rm snd}}$ are generally complex, non-polynomial functions of $k$ with limits: $\lim_{k\rightarrow 0}\omega^{(a)}_{{\rm snd}/{\rm sh}}(k)=0,\quad\lim_{k\rightarrow 0}{\mathfrak{w}}^{(a)}_{{\rm sh}/{\rm snd}}(k)=\text{constant }c^{(a)}_{{\rm sh}/{\rm snd}}\quad\forall a,$ noting that $\omega^{(a)}_{{\rm snd}/{\rm sh}}(k)$ are hydrodynamic modes, while $\mathfrak{w}^{(a)}_{{\rm sh}/{\rm snd}}(k)$ are non-hydrodynamic modes.
@@
$\begin{split}&{(1-i\omega F_{u}-k^{2}R_{u})}\left[\prod_{a=1}^{N^{\rm sh}_{2}}(\omega-{\mathfrak{w}}_{\rm sh}^{a})\right]=C_{\rm sh}\,,\\ &\big{[}(1-i\omega G_{T}-k^{2}R_{T})(1-i\omega F_{u}-k^{2}R_{u})\\ &+k^{2}F_{T}~{}G_{u}\big{]}\left[\prod_{a=1}^{N^{{\rm{\rm snd}}}_{2}}(\omega-{\mathfrak{w}}_{{\rm{\rm snd}}}^{a})\right]=C_{{\rm{\rm snd}}}.\end{split}$ (16)
where $C_{\rm sh}=\lim_{k\rightarrow 0}\left(\prod_{a=1}^{N_{2}^{\rm sh}}\left[-{\mathfrak{w}}^{a}_{\rm sh}\right]\right)$, and $C_{{\rm{\rm snd}}}=\lim_{k\rightarrow 0}\left(\prod_{a=1}^{N_{2}^{{\rm{\rm snd}}}}\left[-{\mathfrak{w}}^{a}_{{\rm{\rm snd}}}\right]\right)$.
@@
If $F_{u(T)}$, $G_{u(T)}$, and $R_{u(T)}$ are finite polynomials in $\omega$ and $k^{2}$, the condition in (16) cannot be fulfilled. However, by permitting an infinite-order derivative expansion in Fourier space (around $\omega=0$ and $k=0$), we can, in principle, eliminate all non-hydrodynamic modes from the dispersion polynomial through an appropriate selection of expansion coefficients $f_{m,n}^{u(T)}$, $g^{u(T)}_{m,n}$, and $R^{u(T)}_{m}$. Moreover, the radius of convergence for this infinite-order frame transformation in Fourier space will be dictated by the position of the nearest non-hydrodynamic mode to the origin in the multi-dimensional space of real momenta and complex frequencies.
Let’s normalize the perturbed solution in the sound channel as: $u_{\mu}^{(s)}=\bar{u}_{\mu}+\epsilon\,\{0,{\vec{k}/k}\}\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$, $T^{(s)}=\bar{T}+\epsilon\,\delta T\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$, where $\delta T$ is dictated by ${\cal R}(\omega,k)$ before the frame transformation and ^$\hat{\cal R}(\omega,k)$ after the transformation. The equivalence of the hydrodynamic sector requires that:
$\begin{split}{\cal R}(\omega^{a}_{{\rm{\rm snd}}},k)=\hat{\cal R}(\omega^{a}_{{\rm{\rm snd}}},k)\quad\forall a=\{1,2,\dots,N_{1}^{{\rm{\rm snd}}}\}\,.\end{split}$ (17)
Let’s examine whether we can obtain a solution for this set of coupled nonlinear equations ((16) and (17)) and to what degree the solution would be unique.

### v2

@@ Abstract
In relativistic hydrodynamics, defining fluid velocity and temperature beyond the perfect fluid case introduces ambiguity. In thermodynamic equilibrium, these variables are naturally defined by uniform thermodynamic potentials. Outside equilibrium, they can be redefined with arbitrary corrections that become nonzero with spacetime variations. These ambiguities are typically resolved by imposing constraints on the conserved currents, akin to gauge fixing in gauge theory, with common choices being the ‘Landau’ or ‘Eckart’ frame[7, 8]. However, one can always opt for constraints beyond these or even proceed without addressing these field ambiguities, as in BDNK theory[9, 10, 11]. It appears that the spectrum of linearized perturbations is significantly altered when the fluid fields are redefined. The BDNK formalism, in particular, leaves the fluid frame unfixed to establish the frame in which the first-order fluid dynamics preserves causality and stability. This method diverges from the established difficulties related to causality and stability in the traditional Landau or Eckart frame[12, 13, 14, 15, 16, 17]. While rotational and translational symmetries are easily addressed in the analysis, gauge freedoms in defining fluid variables add further complexities. Addressing these complexities is crucial for ensuring that the physical constraints derived from linearized perturbation analysis appropriately represent the underlying physics without any artefacts from arbitrary definitions[18, 19, 20, 21, 22, 23]. It is well understood that the linearized spectrum changes with field redefinition, as the perturbations of the redefined fields correspond to different physical quantities than the original fields[24, 25, 26, 27]. However, the causal nature of a specific fluid model should not be contingent upon the choice of variables used for its description.
@@
We, in this work, provided a novel technique to find out which fluid model is physical under arbitrary field redefinition. The method is simple: we, through an all-order field redefinition, eliminate all non-hydrodynamic modes from the fluid model and analyze the structure of the resulting modified energy-momentum tensor. This technique works provided the original stress-energy tensor (the one with non-hydrodynamic modes, before the field redefinition) has a finite number of terms in derivative expansion. We observe that if the new modified stress-energy tensor, after eliminating the non-hydrodynamic modes, incorporates higher derivative terms of the fluid variables that include up to infinite orders, then the fluid theory we started with is physical, whereas, if the modified stress-energy tensor turns out to have a finite number of terms, then the non-hydrodynamic modes of the original theory must emerge from field redefinition artefacts solely and should not be considered physical. Another criterion we find is whether the dispersion polynomial neatly factorizes into hydrodynamic and non-hydrodynamic modes. If the non-hydrodynamic modes do not appear as distinct finite factors, then they are not frame artefacts[29]. In this and a companion study[29], we investigate the impact of field redefinition on the spectrum of linearized perturbations within relativistic fluid dynamics. We have found that the spectrum of hydrodynamic modes (those with frequencies that vanish as spatial momentum approaches zero) remains unchanged by local field redefinition. In contrast, the spectrum of non-hydrodynamic modes (those with frequencies that approach a finite nonzero value as momenta diminish) is influenced by field redefinition. We also show that non-hydrodynamic modes can be removed from the spectrum of linearized perturbations through a suitable, generally all-order field redefinition. Once this is applied, the spectrum will be simplified to include only the hydrodynamic modes of the original fluid model, and the stress-energy tensor becomes ready for analyzing the physical validity of the original non-hydrodynamic modes. We will refer to this new fluid frame as the “hydro frame”. Even if we start with a stress tensor that contains a finite number of terms, transforming to the “hydro frame” typically yields a stress tensor with an infinite number of terms, incorporating arbitrarily higher-order derivatives of the fluid variables. In Ref. [29], we show how the non-hydrodynamic modes of the original theory could control the hydrodynamic expansion’s validity of the new stress-energy tensor.
@@
where $M_{ij}$ is the linearization matrix, and the spectrum is obtained from the zeros of the determinant of $M_{ij}$. If the equation ${\cal E}$ contains a finite number of derivatives, then $\text{Det}[M]$ will be a finite polynomial in $\omega$ and $k$, resulting in a finite number of zeros. Now consider a field redefinition $\Phi_{i}\rightarrow\Psi_{i}=\Phi_{i}+\Delta\Phi_{i}$, where the shift $\Delta\Phi_{i}$ is a nonlinear function of $\Phi_{i}$ and its derivatives. In the context of fluid dynamics, we assume that at equilibrium, $\Phi_{i}$ and $\Psi_{i}$ coincide, meaning $\Delta\Phi_{i}$ vanishes when evaluated at $\Phi_{i}=\bar{\Phi}_{i}$. Under this field redefinition, the equations of motion transform as ${\cal E}(\Phi)\rightarrow\tilde{\cal E}(\Psi)$. While the set of PDEs $\tilde{\cal E}$ can be fully determined from ${\cal E}$ and the transformation $\Delta\Phi$, it will have a distinctly different structure. We can linearize $\tilde{\cal E}$ to obtain the spectrum, similarly to how we did for ${\cal E}$ in (1), resulting in a different linearization matrix $\tilde{M}_{ij}$. Our objective is to establish the relationship between $M_{ij}$ and $\tilde{M}_{ij}$
@@
As $\Delta\Phi$ is vanishing when $\Phi=\bar{\Phi}$, it must be of order $\mathcal{O}(\epsilon)$ or higher for the field configuration in (1). Thus, $\Delta\Phi$ can be represented by the matrix equation up to ${\cal O}(\epsilon^{2})$
@@
$\Rightarrow{\rm Det}[M]={\rm Det}[\tilde{M}]\,{\rm Det}[{\bf 1}+S].$
The zeros of ${\rm Det}[{\bf 1}+S]$ represent new modes in the $\Phi$ frame that are absent in the $\Psi$ frame. These new zero modes arise from perturbations in the $\Phi$ frame, where the additional terms coming from the linearized field redefinitions are exactly canceled by the linearized fluctuations. In the $\Psi$ frame, these new modes indicate no fluctuations, meaning that they are artefacts of the frame transformation with no physical significance. Importantly, depending on the choice of field redefinitions, these artificial modes may appear unstable or acausal, even if the theory in the $\Psi$ frame is entirely valid. However, in the $\Phi$ frame, there is no straightforward way to distinguish these artefacts from genuine physical modes.
@@
$F_{u(T)}$ $\equiv\sum_{m,n}f_{m,n}^{u(T)}\left[\hat{u}\cdot\partial\right]^{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{n}\,,$ (9)
$G_{u(T)}$ $\equiv\sum_{m,n}g^{u(T)}_{m,n}\left[\hat{u}\cdot\partial\right]^{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{n}\,,$
and $R_{u(T)}=R_{u(T)}[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}]$ do not depend on the operator $(\hat{u}\cdot\partial)$. They are included to account for field redefinitions that remain non-zero as $\omega\rightarrow 0$. These can also be expanded in the small $k$ limit as $R_{u(T)}=\sum_{m}r^{u(T)}_{m}\left[\hat{P}^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\right]^{m}$. Here, $f_{m,n}^{u(T)}$, $g^{u(T)}_{m,n}$, and $r^{u(T)}_{m}$ are functions of temperature, and $\hat{P}^{\mu\nu}$ is the projector orthogonal to $\hat{u}^{\mu}$, $\hat{P}^{\mu\nu}=\eta^{\mu\nu}+\hat{u}^{\mu}\hat{u}^{\nu}$. In the context of linearized analysis, the derivatives $(\hat{u}^{\alpha}\partial_{\alpha})$ and $(\hat{P}^{\alpha\beta}\partial_{\alpha})$ commute, so their ordering in the functions don’t matter. For a rotational and translationally invariant equilibrium case, we have $\bar{u}^{\mu}=\{1,0,0,0\},\quad\bar{T}={\rm constant}$, where the perturbed solution is $u_{\mu}^{(s)}=\bar{u}_{\mu}+\epsilon~{}\delta u_{\mu}\,e^{-i\omega t+i\vec{k}\cdot\vec{x}},\quad T^{(s)}=\bar{T}+\epsilon\,\delta T\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$.
@@
where $\Delta u_{k}$ and $\Delta u_{\perp}$ are defined the same way as $\beta_{k}$ and $\beta_{\perp}$. Now we compute ${\rm Det}\left[{\mathbf{1}}+S\right]$ as
${\cal F}(\omega,k^{2})\equiv{\rm Det}[{\mathbf{1}}+S]=(1-i\omega F_{u}-k^{2}R_{u})\times$
$\left[(1-i\omega G_{T}-k^{2}R_{T})(1-i\omega F_{u}-k^{2}R_{u})+k^{2}F_{T}\,G_{u}\right].\quad$ (12)
We have suppressed the arguments of $F_{u(T)}$, $G_{u(T)}$, and $R_{u(T)}$ for simplicity. From Eq. (6), we note that a general field redefinition of the fluid variables introduces an additional factor, $\mathcal{F}$, to the dispersion polynomial which is a polynomial in $\omega$ and $k^{2}$. This generates new modes in the system. Some remarks on the properties of $F_{u(T)}$ and $G_{u(T)}$: If $F_{u}$ and $G_{T}$ are vanishing, and $F_{T}$ and $G_{u}$ do not depend on the operator $(u\cdot\partial)$ or $(-i\omega)$ in Fourier space, then $\mathcal{F}$ will have no zeros for any $\omega$, resulting in no new modes. However, depending on $F_{T}$ and $G_{u}$, $\mathcal{F}$ might have zeros for some real values of $k$. If every term in $F_{u}$ and $G_{T}$ includes at least one factor of $(P^{\mu\nu}\partial_{\mu}\partial_{\nu})$ or $k^{2}$ in Fourier space, the frequency of these new modes will diverge as $k\rightarrow 0$. If either $F_{u}$ or $G_{T}$ has at least one term without $(P^{\mu\nu}\partial_{\mu}\partial_{\nu})$, at least one new mode will appear with a finite, non-zero frequency as $k\rightarrow 0$, resembling a genuine “non-hydrodynamic” mode. None of the zeros of $\mathcal{F}$ will have the form $\lim_{k\rightarrow 0}\omega(k)=0$. Thus, frame redefinitions will not generate new hydrodynamic modes, and redefining fluid variables will not affect the spectrum of hydrodynamic modes in Fourier space.
@@
Now, let’s reverse this process—starting with the ‘Frame-2’ equations (using the hatted fluid variables $\hat{u}^{\mu}$ and $\hat{T}$) and implementing the inverse transformations from (7) and (8) to return to ‘Frame-1’ equation with fluid variables $u^{\mu}$ and $T$. This inversion would naturally eliminate the extra factor $\mathcal{F}$ from the ‘Frame-2’ dispersion polynomial. From the viewpoint of ‘Frame-2’, this effectively removes a non-hydrodynamic mode from the spectrum through field redefinition. The possibility to fully absorb a non-hydrodynamic mode through a field redefinition implies that such a mode may not be physical. A truly physical non-hydrodynamic mode should never be completely erased by a frame transformation.
In this section, we will investigate an “inverse transformation” to remove a mode from the spectrum and analyze if physical non-hydrodynamic modes persist as expected. The dispersion polynomials in ‘Frame-1’ and ‘Frame-2’, indicated as $P(\omega,k^{2})$ and $\hat{P}(\omega,k^{2})$, respectively, are related as
@@
where $\mathcal{F}$ is given in (12), with
$F_{u(T)}$$\equiv$ $\sum_{m,n}f_{m,n}^{u(T)}\left[-i\omega\right]^{m}\left[-k^{2}\right]^{n},$
@@
Due to the isotropy of the background fluid profile and the constitutive relations, the linearized dynamics in the shear and sound sectors will get disentangled. Consequently, both $\hat{P}(\omega,k^{2})$, $P(\omega,k^{2})$ and $\mathcal{F}(\omega,k^{2})$ must factor into contributions from the shear and sound sectors
@@
where $\hat{P}(\omega,k^{2})=\hat{P}_{\rm sh}(\omega,k^{2})\hat{P}_{{\rm snd}}(\omega,k^{2})$, and $P(\omega,k^{2})=P_{\rm sh}(\omega,k^{2})P_{{\rm snd}}(\omega,k^{2})$. Assume that $P_{\rm sh}(\omega,k^{2})$ has $N^{\rm sh}_{1}$ hydrodynamic modes with frequencies $\omega_{\rm sh}^{a}(k)$ and $N^{\rm sh}_{2}$ non-hydrodynamic modes with $\mathfrak{w}_{\rm sh}^{a}(k^{2})$ where, respectively, $a=\{1,2,\ldots,N^{\rm sh}_{1}\}$ and $a=\{1,2,\ldots,N^{\rm sh}_{2}\}$. And in the sound channel, it has $N^{\rm snd}_{1}$ hydrodynamic modes with $\omega_{\rm snd}^{a}(k)$ $\left(a=\{1,2,\ldots,N^{\rm snd}_{1}\}\right)$ and $N^{\rm snd}_{2}$ non-hydrodynamic modes with $\mathfrak{w}_{\rm snd}^{a}(k^{2})$ $\left(a=\{1,2,\ldots,N^{\rm snd}_{2}\}\right)$. Thus, $P_{\rm sh}(\omega,k^{2})$ and $P_{\rm snd}(\omega,k^{2})$ can be factored as $P_{\rm sh}(\omega,k^{2})=\left[\prod_{a=1}^{N^{\rm sh}_{1}}(\omega-\omega_{\rm sh}^{(a)})\right]\left[\prod_{a=1}^{N^{\rm sh}_{2}}(\omega-{\mathfrak{w}}_{\rm sh}^{(a)})\right]$, and $P_{{\rm snd}}(\omega,k^{2})=\left[\prod_{a=1}^{N^{{\rm snd}}_{1}}(\omega-\omega_{{\rm snd}}^{(a)})\right]\left[\prod_{a=1}^{N^{{\rm{\rm snd}}}_{2}}(\omega-{\mathfrak{w}}_{{\rm{\rm snd}}}^{(a)})\right]$. Both $\omega^{(a)}_{{\rm sh}/{\rm snd}}$ and $\mathfrak{w}^{(a)}_{{\rm sh}/{\rm snd}}$ are generally complex, non-polynomial functions of $k$ with limits: $\lim_{k\rightarrow 0}\omega^{(a)}_{{\rm snd}/{\rm sh}}(k)=0,\quad\lim_{k\rightarrow 0}{\mathfrak{w}}^{(a)}_{{\rm sh}/{\rm snd}}(k)=c^{(a)}_{{\rm sh}/{\rm snd}}\,\forall a$.
@@
${(1-i\omega F_{u}-k^{2}R_{u})}\left[\prod_{a=1}^{N^{\rm sh}_{2}}(\omega-{\mathfrak{w}}_{\rm sh}^{a})\right]=C_{\rm sh}\,,$ (16)
$\left[(1-i\omega G_{T}-k^{2}R_{T})(1-i\omega F_{u}-k^{2}R_{u})+k^{2}F_{T}\,G_{u}\right]$
×$\times\left[\prod_{a=1}^{N^{{\rm{\rm snd}}}_{2}}(\omega-{\mathfrak{w}}_{{\rm{\rm snd}}}^{a})\right]=C_{{\rm{\rm snd}}},$
@@
where $C_{\rm sh}=\lim_{k\rightarrow 0}\prod_{a=1}^{N_{2}^{\rm sh}}\left[-{\mathfrak{w}}^{a}_{\rm sh}\right]$ and $C_{{\rm{\rm snd}}}=\lim_{k\rightarrow 0}\prod_{a=1}^{N_{2}^{{\rm{\rm snd}}}}\left[-{\mathfrak{w}}^{a}_{{\rm{\rm snd}}}\right]$. If $F_{u(T)}$, $G_{u(T)}$, and $R_{u(T)}$ are finite polynomials in $\omega$ and $k^{2}$, the condition, (16), cannot be fulfilled. However, by permitting an infinite-order derivative expansion in Fourier space (around $\omega=k=0$), we can, in principle, eliminate all non-hydrodynamic modes from the dispersion polynomial through an appropriate selection of expansion coefficients $f_{m,n}^{u(T)}$, $g^{u(T)}_{m,n}$, $R^{u(T)}_{m}$. Moreover, the radius of convergence for this infinite-order frame transformation in Fourier space will be dictated by the position of the nearest non-hydrodynamic mode to the origin in the multi-dimensional space of real momenta and complex frequencies. Let’s normalize the perturbed solution in the sound channel as: $u_{\mu}^{(s)}=\bar{u}_{\mu}+\epsilon\,\{0,{\vec{k}/k}\}\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$, $T^{(s)}=\bar{T}+\epsilon\,\delta T\,e^{-i\omega t+i\vec{k}\cdot\vec{x}}$, where $\delta T$ is dictated by ${\cal R}(\omega,k)$ before the frame transformation and ^$\hat{\cal R}(\omega,k)$ after the transformation. The equivalence of the hydrodynamic sector requires that:
${\cal R}(\omega^{a}_{{\rm{\rm snd}}},k)=\hat{\cal R}(\omega^{a}_{{\rm{\rm snd}}},k)\quad\forall a=\{1,2,\dots,N_{1}^{{\rm{\rm snd}}}\}\,.$ (17)